OpenStudy (anonymous):

what is the prime factorization of 3000056

OpenStudy (mathmate):

Whenever you have to factorize an even number, find all the factors of two before starting other tests. To give you a kickstart, I'd begin with: \(3000056=2^3*375007\) Now you would have to do divisibility tests for 3,5,7,11,13,17,19,23,29,31,37... (primes) If you don't know the divisibility tests, post. Beyond 13 you would have to divide each time to see if there is any remainder.

OpenStudy (anonymous):

To check your answer http://www.wolframalpha.com/input/?i=factor+3000056

ganeshie8 (ganeshie8):

There is an interesting method to factor huge odd numbers, let me see if i get this correctly.. .we start with : \[375007 = x^2-y^2\]

OpenStudy (anonymous):

try 6064 6033

ganeshie8 (ganeshie8):

Right, but its hard to find them

ganeshie8 (ganeshie8):

\[375007 = x^2-y^2 = (x+y)(x-y)\] you're done once you found the x and y values did u use pell's equation or something to find those values ?

OpenStudy (anonymous):

thats why we have hmm the uncivilized Sieve of Eratosthenes

ganeshie8 (ganeshie8):

this method is called Fermat-Kraitchik factorization supposed to be superior to sieve of eratosthenes

OpenStudy (anonymous):

see for sieve we have testes up to 612 for Fermat i can see 6033 so i guess its a bit miss when u have huge number that have only 2 primes factors other wise it works nice

OpenStudy (mathmate):

Fermat's should work very well for primes for breaking codes, because the factors are usually close together to make huge factors.

ganeshie8 (ganeshie8):

Yeah Fermat-Kraitchik factorization method works nicely when the factors are close together

ganeshie8 (ganeshie8):

but in the present example the number factors to 31*12097 and the method is no longer efficient

OpenStudy (mathmate):

31 is a reasonable number for trial division.